OpenStudy (anonymous):

HELP??? Given:RQ=5 and 2(PQ)+3(RQ)=27 . Use the Substitution Property of Equality to find the value of PQ.

OpenStudy (anonymous):

wheres the diagram?

OpenStudy (anonymous):

no diagram was given, that's why im confused.

jimthompson5910 (jim_thompson5910):

let PQ = x 2(PQ)+3(RQ)=27 2(PQ)+3(5)=27 ... replace RQ with 5 2(x)+3(5)=27 ... replace PQ with x 2x + 15 = 27 now solve for x

OpenStudy (anonymous):

much appreciated!!

jimthompson5910 (jim_thompson5910):

you're welcome